Sherman has budgeted sales for the upcoming quarter as follows:

April May June
Units 1,600 1,900 1,750

The desired ending finished goods inventory for each month is one-half of next month's budgeted sales. Three pounds of direct material are required for each unit produced. If direct material costs $5 per pound, and must be paid for in the month of purchase, the budgeted direct materials purchases (in dollars) for April are:
Answer #1

Sales = 1,600 units

Desired ending inventory = 1,900 / 2 = 950

Total finished units = 1,600 + 950 = 2,550

For 1 unit of finished unit, 3 pounds direct material is required.

Direct material cost is $5 per pound.

For 2,550 finished units the cost is = 2,550 X 3 X $5 = $38,250

Direct materials purchases (in dollars) = $38,250